OpenStudy (anonymous):

wrte the equation of the ellipse with its center at the origin. the height of the ellipse is 12 units and the width is 14 units.

OpenStudy (jdoe0001):

it's wider than it's taller that means horizontal traverse major axis, or is a horizontal ellipse the bigger unit is 14 so half that is our "a" component a = 7 the smaller unit is 12 so half that is our "b" component b = 6 center is at (0, 0) \(\bf \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\)

OpenStudy (anonymous):

X^2/36 + y^2/49=1 would be the answer?

OpenStudy (jdoe0001):

\(\bf \cfrac{x^2}{49}+\cfrac{y^2}{36}=1\)

OpenStudy (jdoe0001):

in an ellipse, the major traverse axis carries the bigger denominator